# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v1

@@ Acknowledgements
This publication is part of the project ‘Clash of the Titans: deciphering the enigmatic role of cosmic collisions’ (with project number VI.Vidi.193.113 of the research programme Vidi which is (partly) financed by the Dutch Research Council (NWO). We thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high-performance computing cluster. We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius.
@@ Appendix A Results as a function of Sérsic index
We analyse how the performance of both methods depends on Sérsic index. In Fig. 12 we compare the predicted $f_{AGN}$ values from Zoobot and GALFIT for galaxies with low Sérsic index ($n<1$) on the top panels and galaxies with high Sérsic index ($n>6$) on the bottom panels. The results from Zoobot for low Sérsic index galaxies have $\text{RMSE}=0.035$, $\langle{\Delta f_{AGN}}\rangle=0.0027$ and $\sigma(\Delta f_{AGN})=0.035$, which corresponds to an increase of 30% in the RMSE. For high Sérsic galaxies, the performance is slightly better than the average for all galaxies, with $\text{RMSE}=0.020$, $\langle{\Delta f_{AGN}}\rangle=0.0004$ and $\sigma(\Delta f_{AGN})=0.021$. On the other hand, low Sérsic index galaxies have, from GALFIT results, a 47% worse RMSE compared to the whole test sample, with $\text{RMSE}=0.133$, $\langle{\Delta f_{AGN}}\rangle=-0.048$ and $\sigma(\Delta f_{AGN})=0.12$, and it appears that the results are worse when not only the Sérsic index is low, but also the $f_{AGN}$ (this is also seen in Fig. 13), where we show how the RMSE varies in different bins of $n$ for the whole sample, but also for galaxies with low and high AGN fraction ($f_{AGN}<0.2$ and $f_{AGN}>0.8$, respectively). A high Sérsic index does not seem to impact the predictions of $f_{AGN}$ determined from GALFIT, with $\text{RMSE}=0.094$, $\langle{\Delta f_{AGN}}\rangle=-0.005$ and $\sigma(\Delta f_{AGN})=0.094$.
Figure 12: Comparison between the real injected AGN fraction and the recovered AGN fraction obtained from the Zoobot (top row) and GALFIT (bottom row), for galaxies with low Sérsic index ($n<1$, left panels) and high Sérsic index ($n>6$, right panels). The colour bar indicates the number of sources in each 2D bin.

### v2

@@ Acknowledgements
This publication is part of the project ‘Clash of the Titans: deciphering the enigmatic role of cosmic collisions’ (with project number VI.Vidi.193.113 of the research programme Vidi which is (partly) financed by the Dutch Research Council (NWO). We thank Mara Salvato for helpful discussions on AGN. We thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high-performance computing cluster. We thank SURF (www.surf.nl) for the support in using the National Supercomputer Snellius.
@@ Appendix A Results as a function of Sérsic index
We analyse how the performance of both methods depends on Sérsic index. In Fig. 15 we compare the predicted $f_{AGN}$ values from Zoobot and GALFIT for galaxies with low Sérsic index ($n<1$) on the top panels and galaxies with high Sérsic index ($n>6$) on the bottom panels. The results from Zoobot for low Sérsic index galaxies have $\text{RMSE}=0.017$, $\langle{\Delta f_{AGN}}\rangle=-0.0025$ and $\sigma(\Delta f_{AGN})=0.017$, which corresponds to an increase of 31% in the RMSE. For high Sérsic galaxies, the performance is slightly better than the average for all galaxies, with $\text{RMSE}=0.013$, $\langle{\Delta f_{AGN}}\rangle=-0.0007$ and $\sigma(\Delta f_{AGN})=0.012$. On the other hand, low Sérsic index galaxies have, from GALFIT results, a 47% worse RMSE compared to the whole test sample, with $\text{RMSE}=0.16$, $\langle{\Delta f_{AGN}}\rangle=-0.073$ and $\sigma(\Delta f_{AGN})=0.14$, and it appears that the results are worse when not only the Sérsic index is low, but also the $f_{AGN}$ (this is also seen in Fig. 16), where we show how the RMSE varies in different bins of $n$ for the whole sample, but also for galaxies with low and high AGN fraction ($f_{AGN}<0.2$ and $f_{AGN}>0.8$, respectively). A high Sérsic index does not seem to impact the predictions of $f_{AGN}$ determined from GALFIT, with $\text{RMSE}=0.094$, $\langle{\Delta f_{AGN}}\rangle=-0.005$ and $\sigma(\Delta f_{AGN})=0.094$.
Figure 15: Comparison between the real injected AGN contribution fraction and the recovered AGN contribution fraction obtained from the Zoobot (top row) and GALFIT (bottom row), for galaxies with low Sérsic index ($n<1$, left panels) and high Sérsic index ($n>6$, right panels). The colour bar indicates the number of sources in each 2D bin.
@@
Appendix B AGN example images
We show in Fig. 17 the 18 galaxies in our sample for which our DL model predicts $f_{\rm AGN}>0.7$. This galaxies display a bright central source. In Fig. 18 and 19 we show random subsets of X-ray and MIR AGN, respectively, with very low $f_{\rm AGN}$ predicted by the model. These galaxies do not show a central component in the F150W filter.
Figure 17: Example galaxies with $f_{AGN}>0.7$. Images are 3.84″across and are displayed with an inverse arcsinh scaling.
Figure 18: Example of X-ray AGN with $f_{AGN}<0.065$. Images are 3.84″across and are displayed with an inverse arcsinh scaling.
Figure 19: Example of MIR AGN with $f_{AGN}<0.065$. Images are 3.84″across and are displayed with an inverse arcsinh scaling.